I want to know how to calculate gross pay if I know the NET amount of my paycheck and the % of deductions.
In this case the NET is $1,056.70 and the deduction amount is 27.6% so what is the GROSS? What is the formula?

Answer by Theo(13342)

the net pay is 1056.7
the deductions from the gross pay are 27.6%.
let g = the gross pay
let n = the net pay
let d = the deduction rate (percent / 100 = rate) = 27.6/100 = .276
the general formula is:
n = g - d * g
factor out the g to get:
n = g * (1 - d)
you are given that n = 1056
you are given that d = .276
the formula becomes:
1056.7 = g * (1 - .276)
simplify to get:
1056.7 = .724 * g
solve for g to get:
g = 1056.7/.724 = 1459.530387.
that's your gross.
working back from this, you know the the net pay is equal to the gross minus .276 * the gross which is equal to 1459.530387 minus .276 * 1459.530387.
that makes the net pay equal to 1056.7.
the solution is confirmed to be good.
the solution is that the gross pay is equal to 1459.530387.
the formula i used is:
n = g * (1 - d)
that translates to:
net pay = gross pay * (1 minus (deduction percent divided by 100)).
solve for gross pay and you get:
gross pay = net pay divided by (1 minus (deduction percent divided by 100)).
that gets you gross pay = 1056.7 / (1 - 27.6/100) = 1459.530387.